OpenStudy (anonymous):

Solve for x:-8(x+8)=-8(x+1)-4 (show the steps to solving this equation)

OpenStudy (da_scienceman):

pls expand the brackets to get ur expression, then evaluate gently

OpenStudy (da_scienceman):

that is step 1 becomes x - 8x-8 = - 8x - 8 -4

OpenStudy (da_scienceman):

pls simplifx further

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Expand each side of the equation.\[-64-8 x=-12-8 x\]Move -8 to the LHS and simplify.\[-52\neq 0 \]There is no solution.

OpenStudy (anonymous):

\[8(x+8)=−8(x+1)−4\] \[=>8x+64=−8x−8−4=−8x−12\] \[=>8x+8x=−64−12=−76\] \[>16x=−76\] \[=>x= \frac{−76}{16}=\frac{19}{4}=4.75..\]